OpenStudy (anonymous):

The table shows data for a class's mid-term and final exams: Mid-Term Final 98 99 98 93 93 93 90 93 88 88 82 86 78 80 75 78 75 78 70 78 68 72 Which data set has the largest standard deviation?

OpenStudy (anonymous):

Mid-term exams Final exams They have the same standard deviation. There is not enough information.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@amistre64 help please?! i think its d

OpenStudy (amistre64):

rule of thumb, range/4

OpenStudy (anonymous):

? @amistre64

OpenStudy (amistre64):

otherwise theres alot of calculations

OpenStudy (amistre64):

theres plenty of information lol so it aint d

OpenStudy (anonymous):

dont get whatcu sayiNG @amistre64

OpenStudy (amistre64):

which one has the larger spread ?

OpenStudy (amistre64):

you dont know what range is? or how to divide by 4?

OpenStudy (anonymous):

Final exams? @amistre64

OpenStudy (amistre64):

the left side has a spread of 30: 68 to 98 the right side has a spread of 27: 72 to 99 the bigger a spread is, the bigger its variance, or deviation is in general 30/4 > 27/4 30(4) > 27(4)

OpenStudy (anonymous):

mid-term exams would be the answer?;D @amistre64

OpenStudy (amistre64):

if thats whats on the left ... sure.

OpenStudy (texaschic101):

first we got to find the mean of both sets of data mid term : (98+98+93+90+88+82+78+75+ 75+70+68) / 11 = 915/11 = 83.18 standard deviation is 10.32 finals : mean is 85.27 standard deviation is 8.19

OpenStudy (amistre64):

we dont "have to"

OpenStudy (anonymous):

Now I'm confused is it Mid-Term Or Finals -_- @amistre64 @texaschic101

OpenStudy (texaschic101):

amistre's way is much easier.....mid terms have bigger standard deviation

OpenStudy (anonymous):

so mid-terms is the answer correct?.. @texaschic101

OpenStudy (texaschic101):

yep